OpenStudy (anonymous):

What is the x-intercept and y-intercept of the graph of 4x - 8y = 32? A) x-intercept = (0, -4); y-intercept = (8, 0) B) x-intercept = (-8, 0); y-intercept = (0, 4) C) x-intercept = (8, 0); y-intercept = (0, -4) D) x-intercept = (0, 4); y-intercept = (-8, 0)

OpenStudy (anonymous):

x intercept is the point where the line intersects the x axis. on the x axis, y =0. so substitute y = 0 in the equation to find the x intercept. similarly substitute x=0 in the equation to find the y-intercept.

OpenStudy (anonymous):

what does it mean when he says "x-intercept" = it means the graph is touching the x-axis so, y=0 hence, substitute 0 (zero) for y in your eqaution and you will find the x-intercep (x,0) what does it mean when he says "y-intercept" = it means the graph is touching the y-axis so, x=0 hence, substitute 0 (zero) for x in your eqaution and you will find the y-intercep (0,y)

OpenStudy (anonymous):

IDK

OpenStudy (anonymous):

substitute y = 0 in your equation and post what you get.

OpenStudy (anonymous):

is it D?

OpenStudy (anonymous):

Why are you guessing the answer? just substitute y = 0 in your equation and see what you get.

OpenStudy (anonymous):

4x - 8y = 32? -8y = 32 - 4x y = 32/(-8) -4x/(-8) y = -4 + (1/2)x this is you equation. Plug in 0 into X and you'll find the Y-intercept (0,y) Plug in 0 into Y and you'll find the X-intercep (x,0)

OpenStudy (anonymous):

A is the answer?

OpenStudy (anonymous):

no, try again.

OpenStudy (anonymous):

C is my final answer lol

OpenStudy (anonymous):

Yup yo got it

OpenStudy (anonymous):

YES!